Circular No. 7536

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000fb IN ESO 205-G11 R. Chassagne, Ste. Clotilde, Île de Réunion, reports his discovery of an

apparent supernova (mag 15.8) on CCD images taken with an automated 0.30-m telescope on Nov. 29.96 and 30.01, and on Dec. 1.50 and 3.86 UT. SN 2000fb is located at α = 5^{h}57^{m}41^{s}.95, δ = −51^{o}57′48′′.0 (equinox 2000.0), which is 6′′ east and 17′′ north of the center of ESO 205-G11. Nothing appears at this position on an image taken by Chassagne on Nov. 10.01 (limiting mag 18) or on the Digital Sky Survey image taken on 1978 Jan. 12 (limiting mag 22).

SUNGRAZING COMETS Further to IAUC 7534, D. Hammer reports his measurements for sev-

eral additional sungrazing comets found by X. Leprette (1997 comets) and by M. Oates (remaining comets) on SOHO website images obtained with the C2 coronagraph. The reduced observations and orbital elements by B. G. Marsden are given on the MPECs cited below.

Comet  1997 UT  α_{2000}  δ_{2000}  MPEC
C/1997 Y2 Dec. 21.336  18^{h}05.^{m}4  −25^{o}01′  2000-X14
C/1997 Y3  22.877  18 12.4  −24 51  2000-X14
Comet  1998 UT  α_{2000}  δ_{2000}  MPEC
C/1998 X3 Dec. 5.538  16^{h}49.^{m}7  −24^{o}12′  2000-X13
C/1998 X4  7.854  17 00.2  −24 29  2000-X14
C/1998 X5  9.188  17 07.3  −24 36  2000-X14
C/1998 X6  9.704  17 09.5  −24 43  2000-X14
C/1998 X7  9.896  17 10.7  −24 44  2000-X14
C/1998 X8  11.896  17 19.6  −24 55  2000-X14
Comet  2000 UT  α_{2000}  δ_{2000}  MPEC
C/2000 V3 Nov. 10.904  14^{h}56.^{m}5  −19^{o}46′  2000-W18

COMET 41P/TUTTLE-GIACOBINI-KRESÁK This comet appears to be in outburst, as indicated by the following

visual m_{1} estimates: Nov. 27.53 UT, 10.2: (A. Hale, Cloudcroft, NM, 0.41m reflector); 28.83, 10.5 (Y. Nagai, Yamanashi, Japan, 0.32-m reflector); Dec. 1.82, 10.4 (M. Tsumura, Wakayama, Japan, 0.32-m reflector).

2000 December 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green